# Kinetically Modified Palatini Inflation Meets ACT Data

### v1

@@ I Introduction
Working in the context of the metric formulation of gravity – and in the reduced Planck units with $m_{\rm P}=M_{\rm P}/\sqrt{8\pi}=1$ – it is well known old ; nmi ; nmiq ; roest that the presence of a non-minimal coupling function
@@
$n_{\rm s}=0.9743\pm 0.0034~{}~{}\mbox{and}~{}~{}r\leq 0.0038.$ (4)
This result significantly affects the viability of various well-known inflationary models plin . As a consequence, several modifications have been recently proposed to reconcile with the data, e.g., (non-)minimal actlinde ; actattr ; aoki ; maity ; rhb ; nmact ; yin ; warm ; rhc ; rha ; act1 ; act3 ; act2 or Starobinsky actellis ; r2drees ; ketov ; acttamv ; r2a ; r2b inflation.
@@ II Set-up
Here $\mathfrak{g}$ is the determinant of the background Friedmann-Robertson-Walker metric, $g^{\mu\nu}$ with signature (+,,,)(+,-,-,-)( + , - , - , - ) and we allow for a kinetic mixing through the function $f_{\rm K}(\phi)$ which is parameterized in term of $f_{\cal R}$ as follows
@@
is an indicative interval for $m$ which assures an ample variation of $f_{\rm K}$. Contrary to the models in Ref. martin – with which ours share the same name – we do not violate the quadratic form of the kinetic operator ˙$\dot{}\phi^{2}$ for $\mu=\nu=0$. By performing a conformal transformation nmi , according to which we define the Einstein frame (EF) metric $\widehat{g}_{\mu\nu}=f_{\cal R}\,g_{\mu\nu}$, we can write ${\sf S}$ in the EF as follows
@@ III Inflationary Dynamics
A period of slow-roll nMI is determined in the EF by the condition:
@@
where the symbol $,\chi$ as subscript denotes derivation w.r.t the field $\chi$. For the expressions above we employ $J$ in Eq. (8), without expressing explicitly $\widehat{V}_{\rm I}$ in terms of $\widehat{\phi}$. For some $m$ values, actually, this is not doable due to the complicate form of the $\widehat{\phi}(\phi)$ function which is the following
@@
For $c_{\cal R}\gg 1$, we are able to solve Eq. (14) w.r.t $\phi_{\star}$ and pursue our analytical approach. Indeed we obtain
@@
It is clear that there is a lower bound on $c_{\rm K}$, above which $\phi_{\star}\leq 1$ and so, our proposal can be stabilized against corrections from quantum gravity. I.e.,
@@ IV Effective Cut-Off Scale
To this end we follow the most restrictive approach analyzing the small-field behavior of our models in the EF – cf. Ref. riotto ; un1 ; un2 ; nmikr . We focus on the second term in the right-hand side of Eq. (7) – or Eq. (27a), see Sec. VI below – for $\mu=\nu=0$ and we expand it about $\langle{\phi}\rangle=0$ in terms of $\widehat{\phi}$ – see Eq. (8) – which coincides with $\phi$ since $\langle{J}\rangle=1$. Our result for some $m$ values of can be written as
@@
independently of $m$. If we reinstall $m_{\rm P}$ in the expressions above, so that we obtain dimensionless quantities in the brackets, and introduce $c_{\cal R}$ in the denominators, we can conclude that the UV cut-off scale is
@@
If we take into account Eq. (17), the condition above yields an upper bound on $c_{\cal R}$ for every selected $n$ and $m$. As we show in Sec. V, there is sizable parameter space consistent with this requirement – even for the most restrictive case with $n=2$. The validity of Eq. (23) is of crucial importance for the viability of the effective theory riotto , independently from the hierarchy between $\phi_{\star}$ and $\Lambda_{\rm UV}$, since it is widely believed that dangerous loop-corrections depend on the energy scale and not on the field values. Therefore, the fact that in our models we obtain $\phi_{\star}\gg\Lambda_{\rm UV}$ – which causes, in principle, concerns regarding corrections from non-renormalizable terms associated with $\Lambda_{\rm UV}$ – does not invalidate our proposal. In addition, a more elaborated derivation of $\Lambda_{\rm UV}$ un0 in a model with $n=4$ and $m=0$ reveal the absence of any relevant problem.
Fig. 1: Allowed curves in the $n_{\rm s}-r$ plane for $n=2$ (left panel) and $4$ (right panel) and various $m$’s indicated on the curves. We take ${N_{\star}}=50$ (dashed lines) or ${N_{\star}}=60$ (solid lines) and minimal (thick lines) or maximal (thin lines) $c_{\cal R}$ values. The marginalized joint $68\%$ [$95\%$] regions from P-ACT-LB-BK18 data are depicted by the dark [light] shaded contours.
@@ V Numerical Results
we can restrict $\lambda$ and $\phi_{\star}$ and compute the model predictions via Eqs. (18a) and (18b), for any selected $c_{\cal R}$, $m$ and $n$ – taking the ${N_{\star}}$ range above we assume a standard post-inflationary evolution plin with high enough reheating temperature – see e.g. Ref. rehpal ; rehpal1 . The $c_{\cal R}$ values can be further constrained by the saturation of Eqs. (16) and (23), from which the first one gives the lowest $c_{\cal R}$ value (for $\phi_{\star}=1$) whereas the second yields the largest one.
@@
In the domain of large $c_{\cal R}$ we assure that the perturbative bound $\lambda=3.5$ is not violated. Decreasing $m$, the needed $c_{\cal R}$ values increase and so Eq. (23) becomes more relevant. Note that the graph for $n=4$ is extended to lower $m$ values. E.g., along the dashed thick lines, with decreasing $m$, we obtain
@@ VI Supergravity Embedding
where $D_{\alpha}W=W_{,z^{\alpha}}+K_{,z^{\alpha}}W$ with $W$ being the superpotential. Therefore the desired SUGRA embedding of our models requires the determination of $K$ and $W$.
@@
It is worth noticing that contrary to the embedding of metric nMI in the SUGRA – cf. Ref. nmikr –, $F_{\cal R}$ and $F_{\cal R}^{*}$ enter two different terms in the right-hand side of Eq. (37) and so they give zero contribution into $K_{\Phi\Phi^{*}}$. Note that for $m=1$, $F_{\rm sh}$ and $F_{\cal R}$ are totally decoupled. If, in addition, we take $n=4$ then $K$ is totaly quadratic. Therefore, this is the most well-motivated case from the point of view of SUGRA. Mild (of order $0.1$) variation of the prefactors of the logarithms is expected to have some impact on the observational results – cf. Ref. nmiq ; jhep – at the cost of some tuning, though, that we avoid in our current investigation.
@@
$1$ real $\widehat{\theta}$ $6+3n^{2}c_{\cal R}\widehat{H}_{\rm I}^{2}/2\phi^{2-n/2}f_{\cal R}^{m+1}$
$2$ real $\widehat{s},\widehat{\bar{s}}$ $6/N_{2}+3n^{2}\widehat{H}_{\rm I}^{2}/2\phi^{2}f_{\cal R}^{m+1}$
@@ VII Conclusions
Prompted by the P-ACT-LB-BK18 data which disfavors the conventional nMI nmi ; roest at more than three standard deviations, we proposed a variant within Palatini gravity which assures compatibility with the current observations.
Namely, working along the lines of Ref. nmikr , we considered a non-canonical kinetic mixing, Eq. (6), – involving the exponent $m$ – in the inflaton sector, apart from the non-minimal coupling to gravity, Eq. (1), which is associated with the potential in Eq. (2). Confining $m$ to the range $(0.87~{}[0.5]-10)$ for $n=2$ [$n=4$] we achieved observational predictions which offer a nice covering of the present data on $n_{\rm s}$ – its compatibility especially with the $n=4$ case for low $m$ values is really impressive, as shown in Fig. 1. Part of the resulting $r$ values there could be accessible in the near future. It is notable, also, that our proposal does not rely on modifications of the post-inflationary cosmic evolution as in Ref. r2drees . Moreover, the enhancement of $n_{\rm s}$ for $n=2$ with $m>0$ is more drastic than that presented in Ref. linear ; actlinde . Our solutions can be attained with subplanckian values of the inflaton, requiring large $c_{\cal R}$’s, and without causing serious problems with the perturbative unitarity since the unitarity-violation scale remains well above the inflationary one. It is gratifying, in addition, that a sizable fraction of the allowed parameter space of our models (with $c_{\cal R}\gg 1$) can be studied analytically.

### v2

@@ I Introduction
Working in the context of the metric formulation of gravity – and in the reduced Planck units with $m_{\rm P}=M_{\rm P}/\sqrt{8\pi}=1$ – one can easily verify old ; nmi ; nmiq ; roest that the presence of a non-minimal coupling function
@@
$n_{\rm s}=0.974\pm 0.0068,~{}a_{\rm s}=0.0062\pm 0.0104~{}~{}\mbox{and}~{}~{}r\leq 0.038,$ (4)
[…]
@@ II Set-up
Here $\mathfrak{g}$ is the determinant of the background Friedmann-Robertson-Walker metric, $g^{\mu\nu}$ with signature (+,,,)(+,-,-,-)( + , - , - , - ) and we allow for a kinetic mixing through the function $f_{\rm K}(\phi)$ which is conveniently parameterized in terms of $f_{\cal R}$ as follows
@@
is an indicative interval for $m$ which assures an ample variation of $f_{\rm K}$. Indeed, any value of $f_{\rm K}$ can be revealed by the parametrization of Eq. (6) selecting conveniently $m=\ln f_{\rm K}/\ln f_{\cal R}$. The considered $m$ margin can be characterized as natural, since the values are of order unity, and allows us to overlaid almost the whole observationally favored region, as we show in Sec. V. Contrary to the models in Ref. martin – with which ours share the same name – we do not violate the quadratic form of the kinetic operator ˙$\dot{}\phi^{2}$ for $\mu=\nu=0$. By performing a conformal transformation nmi , according to which we define the Einstein frame (EF) metric $\widehat{g}_{\mu\nu}=f_{\cal R}\,g_{\mu\nu}$, we can write ${\sf S}$ in the EF as follows
@@ III Inflationary Dynamics
A period of slow-roll nMI is determined in the EF by the condition
@@
with the symbol $,\chi$ as subscript denoting derivation w.r.t the field $\chi$. For the expressions above we employ $J$ in Eq. (8), without expressing explicitly $\widehat{V}_{\rm I}$ in terms of $\widehat{\phi}$. For some $m$ values, actually, this is not doable due to the complicate form of the $\widehat{\phi}(\phi)$ function which is the following
@@
For $c_{\cal R}\gg 1$, we are able to solve Eq. (14) w.r.t $\phi_{\star}$ and pursue our analytical approach. Indeed, we obtain
@@
It is clear that there is a lower bound on $c_{\cal R}$, above which $\phi_{\star}\leq 1$ and so, our proposal can be stabilized against corrections from quantum gravity. I.e.,
@@ IV Effective Cut-Off Scale
To this end we follow the most restrictive approach analyzing the small-field behavior of our models in the EF – cf. Ref. riotto ; un1 ; un2 ; nmikr . We focus on the second term in the right-hand side of Eq. (7) – or Eq. (29a), see Sec. VI below – for $\mu=\nu=0$ and we expand it about $\langle{\phi}\rangle=0$ in terms of $\widehat{\phi}$ – see Eq. (8) – which coincides with $\phi$ since $\langle{J}\rangle=1$. Our result for some $m$ values can be written as
@@
independently of $m$. If we reinstall $m_{\rm P}$ in the expressions above, so that we obtain dimensionless quantities in the bracket, and introduce $c_{\cal R}$ in the denominators, we can conclude that the UV cut-off scale is
@@
If we take into account Eq. (17), the condition above yields an upper bound on $c_{\cal R}$ for every selected $n$ and $m$. As we show in Sec. V, there is sizable parameter space consistent with this requirement – even for the most restrictive case with $n=2$. The validity of Eq. (23) is of crucial importance for the viability of the effective theory riotto , independently from the hierarchy between $\phi_{\star}$ and $\Lambda_{\rm UV}$, since it is widely believed that dangerous loop-corrections depend on the energy scale and not on the field values. Therefore, the fact that in our models we obtain $\phi_{\star}\gg\Lambda_{\rm UV}$ – which causes, in principle, concerns regarding corrections from non-renormalizable terms associated with $\Lambda_{\rm UV}$ – does not invalidate our proposal. In addition, a more elaborated derivation of $\Lambda_{\rm UV}$ un0 for $n=4$ and $m=0$ reveals the absence of any relevant problem.
Fig. 1: Allowed curves in the $n_{\rm s}-r$ plane for $n=2$ (left panel) and $4$ (right panel) and various $m$’s indicated on the curves. We take ${N_{\star}}=50$ (dashed lines) or ${N_{\star}}=60$ (solid lines) and minimal (thick lines) or maximal (thin lines) $c_{\cal R}$ values. The marginalized joint $68\%$ [$95\%$] c.l. regions from P-ACT-LB-BK18 data are depicted by the dark [light] shaded contours.
@@ V Numerical Results
we can restrict $\lambda$ and $\phi_{\star}$ and compute the model predictions via Eqs. (18a) and (18b), for any selected $c_{\cal R}$, $m$ and $n$. Taking the ${N_{\star}}$ range above we assumed that nMI is followed in turn by a oscillatory phase with mean equation-of-state parameter $w_{\rm rh}$, radiation and matter domination, with variation plin ; rehpal ; rehpal1 ; reh1 ; reh2 of $w_{\rm rh}$ and the reheating temperature $T_{\rm rh}$ in the following ranges respectively
$0\lesssim w_{\rm rh}\lesssim 1/3~{}~{}\mbox{and}~{}~{}1~{}{\rm TeV}\lesssim T_{\rm rh}\lesssim 1~{}{\rm EeV},$ (25)
with units reinstalled. The $c_{\cal R}$ values can be further constrained by the saturation of Eqs. (16) and (23), from which the first one gives the lowest $c_{\cal R}$ value (for $\phi_{\star}=1$) whereas the second yields the largest one.
@@
In the domain of large $c_{\cal R}$ we assure that the perturbative bound $\lambda=3.5$ is not violated. Decreasing $m$, the needed $c_{\cal R}$ values increase and so Eq. (23) becomes more relevant. For $n=2$ and $m=0.88$ and $0.87$ the lower and upper bounds on $c_{\cal R}$ from Eqs. (16) and (23) respectively coincide and so the the two (thick and thin) dashed and solid lines intersect. On the other hand, the graph for $n=4$ is extended to lower $m$ values and such an intersection does not occur for the range of $m$ in Eq. (6). E.g., along the dashed thick lines, with decreasing $m$, we obtain
@@
Note, finally, that along the curves depicted in Fig. 1 the obtained $a_{\rm s}$ remains negligibly small and therefore it lies within the margin in Eq. (4). E.g., along the dashed [solid] thick lines, it is $a_{\rm s}\simeq 5.3\cdot 10^{-4}~{}[a_{\rm s}\simeq 3.6\cdot 10^{-4}]$ for $n=2$ whereas for $n=4$, we obtain
$5.2\lesssim-a_{\rm s}/10^{-4}\lesssim 6.3~{}[3.7\lesssim-a_{\rm s}/10^{-4}\lesssim 4.3]$ (28)
with decreasing $m$. These results are of the same order of magnitude with those obtained within conventional (metric or Palatini) nMI.
@@ VI Supergravity Embedding
where $D_{\alpha}W=W_{,z^{\alpha}}+K_{,z^{\alpha}}W$ with $W$ being the superpotential. Therefore, the desired SUGRA embedding of our models requires the determination of $K$ and $W$.
@@
It is worth noticing that contrary to the embedding of metric nMI in SUGRA – cf. Ref. nmikr –, $F_{\cal R}$ and $F_{\cal R}^{*}$ enter two different terms in the right-hand side of Eq. (39) and so they give zero contribution into $K_{\Phi\Phi^{*}}$. Note that for $m=1$, $F_{\rm sh}$ and $F_{\cal R}$ are totally decoupled. If, in addition, we take $n=4$ then $K$ is totaly quadratic. Therefore, this is the most well-motivated case from the point of view of SUGRA. Mild (of order $0.1$) variation of the prefactors of the logarithms is expected to have some impact on the observational results – cf. Ref. nmiq ; jhep – at the cost of some tuning, though, that we avoid in our current investigation.
@@
$1$ real $\widehat{\theta}$ $(6+3n^{2}c_{\cal R}/2\phi^{2-n/2}f_{\cal R}^{m+1})\widehat{H}_{\rm I}^{2}$
$2$ real $\widehat{s},\widehat{\bar{s}}$ $(6/N_{2}+3n^{2}/2\phi^{2}f_{\cal R}^{m+1})\widehat{H}_{\rm I}^{2}$
@@ VII Conclusions
Prompted by the P-ACT-LB-BK18 data actin which is just marginally compatible with the conventional (metric or Palatini) nMI nmi ; roest ; attrpal , we proposed a variant within Palatini gravity which assures comfortable consistency with the current observations.
Namely, working along the lines of Ref. nmikr , we considered a non-canonical kinetic mixing, Eq. (6), – involving the exponent $m$ – in the inflaton sector, apart from the non-minimal coupling to gravity, Eq. (1), which is associated with the potential in Eq. (2). Confining $m$ to the range $(0.87~{}[0.5]-10)$ for $n=2$ [$n=4$] we achieved observational predictions which offer a nice covering of the present data on $n_{\rm s}$ – its compatibility especially with the $n=4$ case for low $m$ values is really impressive, as shown in Fig. 1. Part of the resulting $r$ values there could be accessible in the near future. Our solutions can be attained with subplanckian values of the inflaton, requiring large $c_{\cal R}$’s, and without causing serious problems with the perturbative unitarity since the unitarity-violation scale remains well above the inflationary one. It is gratifying, in addition, that a sizable fraction of the allowed parameter space of our models (with $c_{\cal R}\gg 1$) can be studied analytically.
@@
One possible shortcoming of our proposal is that it employs one extra parameter ($m$) compared to the original model of nMI reducing, thereby, its predictability. As another consequence, however, the enhancement of $n_{\rm s}$ for $n=2$ with $m>0$ may be more drastic than that presented in Ref. linear ; actlinde . Moreover, within our scheme, the reheating phase is not constrained as, e.g., in Ref. rha ; rhc ; act5 ; r2drees ; rhb and corrections to the inflationary potential from other sectors of the theory are not required as, e.g., in Ref. act6 ; oxf ; warm .
Acknowledgments
I would like to thank S. Ketov for interesting discussions.